Two heads of lettuce and three pounds of tomatoes cost $2.85. Three heads of lettuce and two pounds of tomatoes cost $2.90. Find the cost of one head of lettuce and the cost of one pound of tomatoes.(an algebric solution can be accepted.)
Answer by edjones(8007)

Let c=heads of lettuce & t=pounds of tomatoes
2c+3t=2.85
3c+2t=2.9
.
6c+9t=8.55 multiply each side by 3
6c+4t=5.8 multiply each side by 2
----------------subtract
==>5t=2.75
t=.55
.
2c+1.65=2.85
2c=1.20
c=.60
.
Ed